A population consists of the elements 3, 5, 7, 8, and 10. What is the population variance?




Round off your answer to the nearest hundredths.

1
Expert's answer
2022-05-19T09:43:12-0400

Population mean:

"\mu=\cfrac{3+5+7+8+10}{5}=6.6."


Population variance:

"\sigma^2=\sum(x_i-\mu)^2\cdot P(x_i),"

"X-\mu=\\
=\begin{Bmatrix}
 3-6.6,5-6.6,7-6.6,8-6.6,10-6.6
\end{Bmatrix}="

"=\begin{Bmatrix}
-3.6, - 1.6,0.4,1.4,3.4
\end{Bmatrix},"

"\sigma^2=(-3.6)^2\cdot \cfrac{1}{5}+(-1.6)^2\cdot \cfrac{1}{5}+\\
+0.4^2\cdot \cfrac{1}{5}+1.4^2\cdot \cfrac{1}{5}+3.4^2\cdot \cfrac{1}{5}=5.84."